# Preconditioned block encodings for quantum linear systems

### v1

@@ 3.3 Preconditioning by quantum multiplication
Figure 7 shows the resulting values of $\kappa_{s}$ for QSVT from Equation 5. All figures have two reference lines. The blue line labelled ’Original’ is the classically computed condition number after diagonal scaling. Results below this line are better than a classical solver without preconditioning. The orange line labelled ’Diagonal’ is the block encoding of the original matrix. Results below this are better that a quantum solver without preconditioning.
As shown, none of the preconditioners reduce $\kappa_{s}$ below the unpreconditioned value. For SPAI, Figure 7(b), the benefit of increasing infill cannot overcome the increase in subnormalisation factor. All of the SPAI results are worse than not doing preconditioning. TPAI, Figure 7(c), fares a little better due to the lower subnormalisation factors but it, too, offers no benefit over not doing preconditioning. CLAI, Figure 7(d), performs the worst except for the smallest matrices due to the rapid increase in subnormalisation factor.
@@ 3.3.1 Preamplified quantum multiplication
This figure of merit, and not the gate count of the block encoding alone, determines the total cost of matrix inversion and other quantum algorithms (up to logarithmic factors), due to the appearance of the subnormalisation $s$ in $\kappa_{s}$ in the number of queries to the block encoding, see Equation 6.
Here, we introduce the new method of preamplified quantum multiplication. In the spirit of methods constructing preamplified block encodings [26, 22], separate parts of the quantum circuit are amplified individually. Singular value amplification allows to improve the subnormalisation of a block encoding by a factor of $\gamma$ with approximately
@@
which can be achieved if the original subnormalisations $\alpha,\beta$ allow for strong amplification. For the example matrices and matrix sizes considered in this work, we find that for preamplification of $U_{A}$, $\gamma_{1}\simeq 1.25$ is the best that can be achieved. For $U_{P}$, $\gamma_{2}\simeq 1.25-2.0$ can be achieved with more levels of infill allowing higher values of $\gamma_{2}$. However, these are not sufficient to reduce the figure of merit, and we use regular quantum multiplication [26]. However, in larger systems or if the matrices involved have worse subnormalisations, our preamplified multiplication method will provide an advantage.
@@ 3.4 Preconditioning by classical multiplication
The second approach is to perform the product of $P$ and $A$ as a classical preprocessing step and the encode $PA$. This is only feasible for TPAI and SPAI as $PA$ retains some degree of sparsity. Figure 8 shows the resulting values of $s$ and $\kappa_{s}$ for encoding $PA$.
@@
Figure 8: Preconditioning by classical multiplication - subnormalisation factors and condition numbers for the product $PA$ using the TPAI and SPAI preconditioners. The blue and orange lines in Figures (c) and (d) are the non-preconditioned reference lines as in Figure 7. Numbers in brackets indicate levels of infill.
The subnormalisation factors, Figure 8(a) and Figure 8(b), for TPAI and SPAI have similar values and are relatively insensitive to the size of the mesh and the number of levels of infill. When encoding the preconditioner, $P$, by itself, the subnormalisation factors for SPAI were significantly larger than TPAI due to the fact that the infill created more diagonals, see Table 1. However, when encoding $PA$, SPAI creates more diagonals than TPAI, but many of the diagonals contain only zeros to within the precision of the matrix product as discussed in Section 2.6. The presence of the zero-valued diagonals in SPAI does not influence the subnormalisation factor as the prep operator just loads a zero coefficient. However, removing them from the encoding reduces the number of qubits needed for the prep register. The net result is that SPAI with 3 levels of infill reduces the QSVT condition number, $\kappa_{s}$, by almost an order of magnitude. This, in turn, leads to a reduction of over 20 in the number of QSVT phase factors.
@@
The classical product $PA$ can be parallelised and since each row-column multiplication leads to one multi-controlled rotation, the encoding circuit can be written directly without the need to form an intermediate matrix. Since the rotations for each diagonal commute, the parallel circuit generation does not create any loop-order dependencies.
Reducing the depth of the encoding circuit is described in the next section and relies on the fact that, as well as creating zero valued diagonals, SPAI also tends to equalise the values along each of the non-zero diagonals.
@@ 4 Query oracle circuit trimming
The circuit trimming approach reduces the circuit size required for block encoding a band diagonal matrix (see Section 2.4. It relies on finding multiplexed rotations where the rotation angles are the same and the multiplexed controls differ by a Hamming distance of 1. The effectiveness of this is enhanced if the matrix is filtered to increase the number of equal-valued entries along each diagonal. We take a digitisation approach where values that are close to each other are placed in bins of equal value. However, the filtering needs to take account of the physical characteristics of the matrix and, for example, just because an entry is close to zero does not mean it can be rounded to zero. We use the double-pass filter approach described in Appendix E. The filter involves a free parameter, $f$ that controls the size of the bins. This is a relative factor that sets the bin size to $\left[(1-\frac{f}{2})\bar{d_{i}},(1+\frac{f}{2})\bar{d_{i}}\right]$ where $\bar{d_{i}}$ is the average value of the $i^{th}$ bin. The first pass creates a list of overlapping bins that all meet the size criterion. The second pass selects non-overlapping bins giving preference to bins with a large number of entries.
@@
Figure 9: Effect of the smoothing parameter on the encoding circuit for $PA$ where $P$ uses SPAI with 3 levels of infill. The numbers of angles and rotations are shown as percentages of the unsmoothed values. Legends refer to the matrix dimensions not the CFD mesh dimensions.
Figure 9 shows the influence of the smoothing factor, $f$, on the encoding of $PA$ for the 16x16, 32x32 and 64x64 CFD meshes. The preconditioner is SPAI with 3 levels of infill. The overall trends are as expected: increasing $f$ reduces the number of unique angles (i.e. fewer bin with more entries in each bin) and rotation gates, but increases the $L_{2}$ error of the solution vector.
There are two other notable features in Figure 9. The first is that the reduction in the number of unique angles is far greater than the reduction in the number of rotation gates. This is because not all multiplexed controls meet the requirement to be coalesced. The second feature is that the percentage reduction in the number of unique values and rotations increases with the size of the mesh. Taking the circuits for which the $L_{2}$ error is $\leq 10^{-2}$, the number of rotations relative to the unsmoothed case is 70.7%, 49.5%, 30.7% for the 16x16, 32x32 and 64x64 meshes respectively. These features are discussed in Section E.3.
@@ 5 CFD Applications using QSVT
The first case is the pressure correction matrix for a 2D cavity with a 32x32 mesh sampled after 100 outer non-linear iterations. Figure 10 compares the QSVT and classical state vector solutions. These are stored at every point in the 2D mesh and are ordered in sweeps across the horizontal grid lines starting at the lower left corner of the cavity and ending at the top right. The results show good agreement with small differences due to the application of oracle trimming. The details of the QSVT calculation are:
@@
  • Subnormalisation factor from classical multiplication, $PA$ $=4.81$
@@
  • $L_{2}$ norm of difference with the classical solution $=2.22\times 10^{-2}$
  • Total number qubits is 17 of which 16 are used for encoding $PA$, (with reference to Figure 2, there are 10 |$\ket{j}$ qubits, 5 |$\ket{s}$ qubits and 1 |$\ket{d}$ qubit) and 1 for signal processing.
As shown in Figure 9 the large reduction in the number of unique values does not translate into an equivalent reduction in the number of rotation gates. For comparison, previous HHL investigations estimated that this case would require 29 qubits [29]. Without preconditioning, QVST would require of the order 350,000 phase factors, a reduction of a factor of 25. Without preconditioning the number of rotation gates to block encode $A$ is 4,290. Hence, the overall circuit depth is reduced by a factor of 12.5 with preconditioning.
@@
Figure 11 compares the classical and QSVT solutions at a cross-section midway along the channel. There is very good agreement between the classical and QSVT solutions. The details of the QSVT calculation are:
@@
  • Subnormalisation factor from classical multiplication, $PA$ $=5.48$
@@
  • $L_{2}$ norm of difference with the classical solution $=8.2\times 10^{-3}$
  • Total number qubits is 20 of which 19 are used for encoding $PA$, (with reference to Figure 2, there are 12 |$\ket{j}$ qubits, 6 |$\ket{s}$ qubits and 1 |$\ket{d}$ qubit) and 1 for signal processing.
For comparison, 879 QSVT phase factors are needed to solve the system without preconditioning. However, the encoding of $A$ requires only 8,348 rotation gates. Hence, in this case, the factor of 3.5 increase in phase factors is more than offset by the 4.9x saving in the number of rotation gates. The larger saving in rotation gates is due to the fact that the Laplacian matrix has a regular structure with many repeated entries, even before smoothing. This is not representative of 3D CFD testcases and savings similar to the 2D testcase are expected.
@@ 6 Conclusions
Although this work has exploited the diagonal matrix structure of lattice-based meshes, for unstructured meshes the benefits of SPAI with infill are expected to be repeated.
@@ 7 Acknowledgements
The permission of Rolls-Royce and Riverlane to publish this work is gratefully acknowledged.
Appendix A Diagonal preconditioning
Diagonal preconditioning is based on scaling the rows and columns of a matrix by the inverse of the diagonal. If $D$ is the matrix formed by the diagonal entries in $A$, then preconditioning is applied by:
$\left(D^{-1}A\right)\ket{x}=D^{-1}\ket{b}$ (11)
@@ Appendix A Diagonal preconditioning
$\left(D_{s}^{-1}AD_{s}^{-1}\right)D_{s}\ket{x}=D_{s}^{-1}\ket{b}$ (12)
The difference between the two formulations is that Equation 11 scales the rows of $A$ and Equation 12 scales both the rows and columns. The former is preferred as the solution vector, |$\ket{x}$, is left unchanged. Classically, this is also referred to as Jacobi preconditioning but is rarely used as there are better performing preconditioners which yield a much greater reduction in the condition number. However, for QLES, the benefit is that as well as setting all the diagonal entries to unity, diagonal scaling tends to equalise other entries in the matrix. This is a necessary step prior to the Toeplitz and circulant approximate inverses discussed in Appendix C and Appendix D.
Note that diagonal preconditioning is not the same as diagonal encoding.
@@ Appendix B Sparse Approximate Inverse (SPAI)
$||A||_{F}=\sqrt{\sum_{i,j}|a_{i,j}|^{2}}=\sqrt{\text{Tr}(AA^{\dagger})}$ (14)
@@ B.1 Sparse-Sparse Iteration Method
$\begin{array}[]{rcl}G_{k}&=&I-AM_{k}\\ \alpha_{k}&=&\text{tr}(G_{k}^{\dagger}AG_{K})/||AG_{k}||_{F}^{2}\\ M_{k+1}&=&M_{k}+\alpha_{k}G_{k}\end{array}$ (15)
Each iteration of the algorithm adds new non-zero entries and reduces the sparsity of $M_{k}$ as shown in Figure 12. This is called infill and the degree of infill is controlled via numerical dropping, i.e. ignoring any entries that fall outside the infill sparsity pattern. This can be applied to either $M_{k}$ or $G_{k}$ [32]. A common choice is to retain the same sparsity pattern for $M_{k}$ as $A_{k}$.
@@
$\alpha_{0}=\frac{||A||_{F}}{||AA^{\dagger}||_{F}}$ (16)
Figure 12: Zoomed in sparsity patterns for SPAI applied to 16x16 pressure correction matrix with 1 level of infill. Left: original matrix ($A$), centre: preconditioning matrix ($P$), right: result of preconditioning ($PA$)
@@ B.2 Column Minimisation Method
Equation 19 can be achieved by a permutation operator, $P_{j}$ that can be applied to $A$ to give:
@@
The size of $s$ depends on the level of infill. For small values of $s$, Equation 22 can be directly inverted. For larger values, it can be solved via a least squares minimisation [12] or by finding the minimum of an energy functional [34]. The key to performance is that the $N$ equations are independent and can be solved in parallel. Here, we use a direct Lower-Upper decomposition to solve Equation 22 exactly to within machine precision.
The infill pattern for the column minimisation method is the same as the sparse-sparse method and is computed by taking successive products with itself. This is an expensive task but can be performed as a one-off calculation. Since only the sparisty pattern is needed, it is sufficient to know whether each row-column multiplication has a common entry.

### v3

@@ 3.3 Preconditioning by quantum multiplication
Figure˜7 shows the resulting values of $\kappa_{s}$ for QSVT from Equation˜5. All figures have two reference lines. The blue line labelled ’Classical’ is the classically computed condition number after diagonal pre-scaling. Results below this line are better than a classical solver without preconditioning. The green line labelled ’No Precon’ is the block encoding of the original matrix. Results below this are better that a quantum solver without preconditioning.
As shown, none of the preconditioners reduce $\kappa_{s}$ below the unpreconditioned value. For SPAI, Figure˜7(b), the benefit of increasing infill cannot overcome the increase in subnormalisation factor. All of the SPAI results are worse than not doing preconditioning. TPAI, Figure˜7(c), fares a little better due to the lower subnormalisation factors but it, too, offers no benefit over not doing preconditioning. CLAI, Figure˜7(d), performs the worst except for the smallest matrices due to the rapid increase in subnormalisation factor.
@@ 3.3.1 Preamplified quantum multiplication
This figure of merit, and not the gate count of the block encoding alone, determines the total cost of matrix inversion and other quantum algorithms (up to logarithmic factors), due to the appearance of the subnormalisation $s$ in $\kappa_{s}$ in the number of queries to the block encoding, see Equation˜6.
Here, we study the method of preamplified quantum multiplication [30]. In the spirit of methods constructing preamplified block encodings [29, 25], separate parts of the quantum circuit are amplified individually to reduce the total figure of merit. Singular value amplification allows to improve the subnormalisation of a block encoding by a factor of $\gamma$ with approximately
@@
which can be achieved if the original subnormalisations $\alpha,\beta$ allow for strong amplification. For the example matrices and matrix sizes considered in this work, we find that for preamplification of $U_{A}$, $\gamma_{1}\simeq 1.25$ is the best that can be achieved. For $U_{P}$, $\gamma_{2}\simeq 1.25-2.0$ can be achieved with more levels of infill allowing higher values of $\gamma_{2}$. However, these are not sufficient to reduce the figure of merit, and we use regular quantum multiplication [29]. However, if the matrices involved have worse subnormalisations, the preamplified quantum multiplication method will provide an advantage.
@@ 3.4 Preconditioning by classical multiplication
The second approach is to perform the product of $P$ and $A$ as a classical preprocessing step and the encode $PA$. This is only feasible for TPAI and SPAI as $PA$ retains some degree of sparsity. Figure˜8 shows the resulting values of $s$ and $\kappa_{s}$ for encoding $PA$.
@@
Figure 8: Preconditioning by classical multiplication - subnormalisation factors and condition numbers for the product $PA$ using the TPAI and SPAI preconditioners. The blue and green lines in Figures (c) and (d) are the non-preconditioned reference lines as in Figure˜7. Numbers in brackets indicate levels of infill.
The subnormalisation factors, Figure˜8(a) and Figure˜8(b), for TPAI and SPAI have similar values and are relatively insensitive to the size of the mesh and the number of levels of infill. When encoding the preconditioner, $P$, by itself, the subnormalisation factors for SPAI were significantly larger than TPAI due to the fact that the infill created more diagonals, see Table 1. However, when encoding $PA$, SPAI creates more diagonals than TPAI, but many of the diagonals contain only zeros to within the precision of the matrix product as discussed in Section˜2.6. The presence of the zero-valued diagonals in SPAI does not influence the subnormalisation factor as the prep operator just loads a zero coefficient. However, removing them from the encoding reduces the number of qubits needed for the prep register. The net result is that SPAI with 3 levels of infill reduces the QSVT condition number, $\kappa_{s}$, by almost an order of magnitude. This, in turn, leads to a reduction of over 20 in the number of QSVT phase factors.
@@
3.4.1 Classical preprocessing considerations
Here, we consider the preprocessing overheads relative to a classical iterative solver of the Krylov subspace family. For the latter, preconditioning involves an extra matrix-vector multiplication on every step. This is an additional prefactor that doesn’t affect the asymptotic complexity $O(Ns\sqrt{\kappa}\log(1/\epsilon))$. For QSVT, preconditioning involves forming the product $PA$ once at the start of the circuit. Multiplying two banded diagonal matrices with $d_{1}$ and $d_{2}$ diagonals requires $O(Nd_{1}d_{2})$ steps. Since this is independent of $\kappa$, which is $O(N^{2/d})$ for $d$-dimensions, it has a lower order complexity an additional matrix-vector multiplication every iteration.
A further consideration is the cost of finding the sparsity patterns for $P$ and $PA$. Both are directly related to the difference stencil used to discretise the flow equations and can be computed, for any level of infill, without the need for matrix-matrix multiplications. Referring to Table˜1, the diagonals with all zero entries in $PA$ can also be determined, a priori, from the difference stencil as described in Section˜B.2. Thus reducing the number of diagonals to be processed and, hence, the scaling prefactor. Since the number of diagonals in $PA$ scales linearly with the number of infill levels, the overhead of needing more infill levels to match the classical condition number (i.e. $\kappa_{s}=\kappa$) is marginal.
SPAI can be easily parallelised since the $N$ systems in Equation˜22 are independent. Logically, we can think of each processor as writing a tape of values which then needs to be used to populate the entries in a sparse matrix. For quantum solvers, the tapes can be used to directly create the encoding circuit without the need to form an intermediate matrix. Both operations are $O(N)$ but the latter has a lower prefactor.
The above complexity arguments extend to 3-dimensional lattice based meshes which cover a large class of CFD applications. They also extended to more general unstructured meshes. However, the compactness of the banded diagonal block encoding is lost and the challenge for unstructured meshes is to find encodings with $O(1)$ subnormalisation factors, as in Figure˜8(b).
@@ 4 Query oracle circuit trimming
The circuit trimming approach reduces the circuit size required for block encoding a band diagonal matrix (see Section 2.4. It relies on finding multiplexed rotations where the rotation angles are the same and the multiplexed controls differ by a Hamming distance of 1. The effectiveness of this is enhanced if the matrix is filtered to increase the number of equal-valued entries along each diagonal. We take a digitisation approach where values that are close to each other are placed in bins of equal value. However, the filtering needs to take account of the physical characteristics of the matrix and, for example, just because an entry is close to zero does not mean it can be rounded to zero. We use the double-pass filter approach described in Appendix˜E. The filter involves a free parameter, $f$ that controls the size of the bins. This is a relative factor that sets the bin size to $\left[(1-\frac{f}{2})\bar{d_{i}},(1+\frac{f}{2})\bar{d_{i}}\right]$ where $\bar{d_{i}}$ is the average value of the $i^{th}$ bin. The first pass creates a list of overlapping bins that all meet the size criterion. The second pass selects non-overlapping bins giving preference to bins with a large number of entries.
@@
Figure 9: Effect of the smoothing parameter on the encoding circuit for PAPA where $P$ uses SPAI with 3 levels of infill. The numbers of angles and rotations are shown as percentages of the unsmoothed values. Legends refer to the matrix dimensions not the CFD mesh dimensions.
Figure˜9 shows the influence of the smoothing factor, $f$, on the encoding of $PA$ for the 16x16, 32x32 and 64x64 CFD meshes. The preconditioner is SPAI with 3 levels of infill. The overall trends are as expected: increasing $f$ reduces the number of unique angles (i.e. fewer bin with more entries in each bin) and rotation gates, but increases the $L_{2}$ error of the solution vector.
There are two other notable features in Figure˜9. The first is that the reduction in the number of unique angles is far greater than the reduction in the number of rotation gates. This is because not all multiplexed controls meet the requirement to be coalesced. The second feature is that the percentage reduction in the number of unique values and rotations increases with the size of the mesh. Taking the circuits for which the $L_{2}$ error is $\leq 10^{-2}$, the number of rotations relative to the unsmoothed case is 70.7%, 49.5%, 30.7% for the 16x16, 32x32 and 64x64 meshes respectively. These features are discussed in Section˜E.3.
@@ 5 CFD Applications using QSVT
The first case is the pressure correction matrix for a 2D cavity with a 32x32 mesh sampled after 100 outer non-linear iterations. Figure˜10 compares the QSVT and classical state vector solutions. These are stored at every point in the 2D mesh and are ordered in sweeps across the horizontal grid lines starting at the lower left corner of the cavity and ending at the top right. The results show good agreement with small differences due to the application of oracle trimming. The details of the QSVT calculation are:
@@
  • Subnormalisation factor from classical multiplication, PAPA =4.81=4.81
@@
  • $L_{2}$ norm of difference with the classical solution $2.22\times 10^{-2}$
[…]
As shown in Figure˜9 the large reduction in the number of unique values does not translate into an equivalent reduction in the number of rotation gates. For comparison, previous HHL investigations estimated that this case would require 29 qubits [33]. Without preconditioning, QSVT would require of the order 350,000 phase factors, a reduction of a factor of 25. Without preconditioning the number of rotation gates to block encode $A$ is 4,290. Hence, the overall circuit depth is reduced by a factor of 12.5 with preconditioning.
@@
Figure˜11 compares the classical and QSVT solutions at a cross-section midway along the channel. There is very good agreement between the classical and QSVT solutions. The details of the QSVT calculation are:
@@
  • Subnormalisation factor from classical multiplication, PAPA =5.48=5.48
@@
  • $L_{2}$ norm of difference with the classical solution $8.2\times 10^{-3}$
[…]
For comparison, 879 QSVT phase factors are needed to solve the system without preconditioning. However, the encoding of $A$ requires only 8,348 rotation gates. Hence, in this case, the factor of 3.5 increase in phase factors is more than offset by the factor 4.9 saving in the number of rotation gates. The larger saving in rotation gates is due to the fact that the Laplacian matrix has a regular structure with many repeated entries, even before smoothing. This is not representative of 3D CFD testcases and savings similar to the 2D testcase are expected.
@@ 7 Acknowledgements
The permission of Rolls-Royce plc and Riverlane to publish this work is gratefully acknowledged.
Appendix A Diagonal Scaling
Diagonal scaling is based on scaling the rows and columns of a matrix by the inverse of the diagonal. Whilst diagonal scaling can be cast as a preconditioner, we use the term scaling to separate it from the other preconditioners. As described in Section˜2.2, diagonal scaling is used to preprocess $A$ prior to encoding for all the preconditioners.
If $D$ is the matrix formed by the diagonal entries in $A$, then preconditioning is applied by:
[…]
@@ Appendix A Diagonal Scaling
[…]
Note that diagonal scaling is not the same as diagonal encoding.
@@ Appendix B Sparse Approximate Inverse (SPAI)
$A||_{F}=\sqrt{\sum_{i,j}|a_{i,j}|^{2}}=\sqrt{\text{Tr}(AA^{\dagger})}$ (14)
@@ B.1 Sparse-Sparse Iteration Method
[…]
Each iteration of the algorithm adds new non-zero entries and reduces the sparsity of $M_{k}$ as shown in Figure˜12. This is called infill and the degree of infill is controlled via numerical dropping, i.e. ignoring any entries that fall outside the infill sparsity pattern. This can be applied to either $M_{k}$ or $G_{k}$ [36]. A common choice is to retain the same sparsity pattern for $M_{k}$ as $A_{k}$.
@@
[…]
Figure 12: Zoomed in sparsity patterns for SPAI applied to 16x16 pressure correction matrix with 1 level of infill. Left: original matrix (AA), centre: preconditioning matrix ($P$), right: result of preconditioning (PAPA)
@@ B.2 Column Minimisation Method
Equation˜19 can be achieved by a permutation operator, $P_{j}$ that can be applied to $A$ to give:
@@
The size of $s$ depends on the level of infill. For small values of $s$, Equation˜22 can be directly inverted. For larger values, it can be solved via a least squares minimisation [12] or by finding the minimum of an energy functional [38]. The key to performance is that the $N$ equations are independent and can be solved in parallel. Here, we use a direct Lower-Upper decomposition to solve Equation˜22 exactly to within machine precision.
The infill pattern for the column minimisation method is the same as the sparse-sparse method. This can be computed by taking successive products of $A$ with itself. However, this is an expensive task which does not exploit any fore knowledge of the matrix. Section˜3.4.1 discusses how the difference stencil of the CFD discretisation can be used to directly compute the infill sparsity pattern. This is illustrated in Figure˜13 for one level of infill. As can be seen, the infill extends the stencil by one grid node in each direction. The product of the matrix and preconditioner further extends the stencil by one more node. Since each SPAI system is solved exactly, the interior nodes in the stencil (Figure˜13(c)) are zero to machine precision. This illustration corresponds to the second row of Table˜1 where the number of non-zero diagonals reduces from 25 to 13.
@@
(a) CFD Matrix, AA.
(b) Preconditioner, $P$, with infill=1.
(c) Product PAPA.
Figure 13: Difference stencil for a single node in the mesh. Solid circles denote non-zero entries in the stencil, open circles denote entries in the stencil for PAPA that are zero by construction. The highlighted points correspond to the entries on a single row of the matrices shown in Figure˜12.